# The Self-Duality Equations on a Riemann Surface and Four-Dimensional Chern–Simons Theory

### v1

@@ 3.1 $\mathbb{CP}^{1}$-Family of 4d CS Theories and the Corresponding 2d Actions
Since there are a whole $\mathbb{CP}^{1}$ family of hyperkähler structures on $\mathscr{M}_{\text{H}}(\Sigma,G)$, it is expected from our computations in §2.4 that there exists a family of 4d Chern–Simons theories that produces this hyperkähler family of symplectic structures. For this matter, we propose the following $\mathbb{CP}^{1}$-family of meromorphic $(1,0)$-form on $C$
@@
giving the meromorphic $(1,0)$-form used in §2.3 that led to the 4d Chern–Simons theory formulation of Hitchin’s equations with symplectic structure $\Omega_{\mathcal{I}}$. We will see that the $\zeta$ above will be the twistor parameter $\zeta$ of the Hitchin hyperkähler structure, parametrizing the symplectic structures in (1.4). Here we will work with complexified fields, Euclidean reality conditions can be imposed in direct analogy with section 2.3.
@@
These boundary conditions coalesce to those in (2.16) as $\zeta\to 0$. We will find it useful to write $\widehat{\zeta}=-1/\bar{\zeta}$ for the stereographic co-ordinate of the point antipodal to $\zeta$.
@@
To determine the resulting 2d action we use the same procedure as explained in the proof of Lemma 2.2, which yields $\frac{\mathrm{i}}{4\pi}S_{\mathrm{H},\zeta}$ where
@@
$-2\mathrm{i}\zeta\mathop{\text{\Large$\int_{\text{\normalsize$\scriptstyle\Sigma$}}$}}\nolimits\mathrm{tr}(k^{-1}\partial k\,\bar{\partial}\widetilde{\eta})-2\mathrm{i}\bar{\zeta}\int_{\Sigma}\mathrm{tr}(\partial\eta\,\bar{\partial}kk^{-1})\,$,
where $S_{\text{WZW}}[k]$ is the Wess–Zumino–Witten action given in equation (2.8).
@@
By taking linear combinations of the above it’s straightforward to see that
@@
The 4d Chern–Simons reduction to Harmonic maps or principal chiral models are treated in detail in [CostelloYamazaki201908]. We first remark that this also uses an $\omega$ of the form (3.1), but the boundary conditions used are to instead impose $\mathsf{A}=0$ at $\zeta$ and $1/\bar{\zeta}$ instead of (3.4). The Wilson line from $\zeta$ to $1/\bar{\zeta}$ then gives the (twisted) harmonic maps of [Donaldson198707, Corlette198801]. The Lax pair thereby obtained are gauge equivalent to ours but the twisted harmonic map is realized in a different gauge from our $k$.
@@ 3.2 Hyperkähler Family of Symplectic Structures
In §3.1, we derive a family of 2d actions (3.10) parametrized by $\zeta\in\mathbb{CP}^{1}$ whose equations of motion, for each value of $\zeta$, are the Hitchin’s equations (1.3). However, our aim was to find the relationship between $\zeta$ and the twistor parameter of the Hitchin moduli space $\zeta$, which determines the family of symplectic structures (1.4). To this aim, we proceed as we did in §2.4 and compute the symplectic structure that 4d Chern–Simons theory with the choice (3.1) associates to $\Sigma\times S^{1}$, as discussed in §2.4.
@@
By varying $\zeta$ over $\mathbb{CP}^{1}$, (3.18) generates the whole $\mathbb{CP}^{1}$-family of symplectic structures associated to the hyperkähler structure on $\mathscr{M}_{\text{H}}(\Sigma,G)$. Our $\mathbb{CP}^{1}$-family of 4d Chern–Simons theory using $\omega(\zeta,\bar{\zeta})$ of (3.1) generates the $\mathbb{CP}^{1}$-family of 2d actions (3.10) giving the symplectic structures $\Omega_{\mathcal{J}_{\zeta}}$ on $\mathscr{M}_{\text{H}}(\Sigma,G)$, these being the Kahler 2-forms for the complex structures $\mathcal{J}_{\zeta}$ on the space of solutions.
@@ 4 Affine Toda Theory
The complexified Hitchin equations are, up to gauge, the generic reduction of the self-dual Yang–Mills from 4d to 2d along a non-degenerate pair of translations. Such reductions are described in [MasonWoodhouse199605, §6.2] and include a number of familiar 2d integrable systems, including harmonic maps, Toda field theory and the sine-Gordon equation. These are distinguished as reductions by gauge choices and choices of ‘constants of integration’ that naturally arise in the reductions. When these models are reformulated as 4d Chern–Simons theories, these choices are encoded in the choice of $\omega$ and the boundary conditions on $\mathsf{A}$. It turns out that the harmonic map reductions require different such choices—these are already well-understood in [CostelloYamazaki201908] and require $\omega$ to have two double poles and two single zeroes with $\mathsf{A}$ vanishing at the double poles and allowed to have poles at the zeroes. However here we see that Toda field theories fit into the choices made for the Hitchin equations in §2 and can be understood as arising from the imposition of further symmetries. This is a continuous symmetry for the Toda chain, i.e., Toda theory for a finite-dimensional simple Lie algebra, but for the affine Toda case we impose a discrete symmetry. In the following we distinguish the Toda chain from affine (also known as extended) Toda.
Our construction is closely related to, but differs from that appearing in [FukushimaSakamotoYoshida202112], where affine Toda theory is engineered using the same $\omega$, but with order defects. We anticipate that our treatment can be connected to theirs by integrating out the local degrees of freedom on the defects.
@@ 4.1 Toda Field Theories
Toda theories can be defined for any Lie algebra but for simplicity we restrict to the case of $\mathfrak{su}(n)$ for the $A_{r-1}$ series and its extended or affine version $\widehat{\mathfrak{su}}(n)$. Both cases can be treated together if we introduce variables $v_{1},\ldots,v_{r}$ with $\sum v_{i}=0$. We then have Lagrangians
$\widehat{L}=\sum_{i=1}^{n}\frac{1}{2}\partial_{w}v_{i}\partial_{\bar{w}}v_{i}-\lambda\mathrm{e}^{v_{i+1}-v_{i}}\,,\hskip 18.49988ptL=\sum_{i=1}^{n}\frac{1}{2}\partial_{w}v_{i}\partial_{\bar{w}}v_{i}-\lambda\sum_{i=1}^{n-1}\mathrm{e}^{v_{i+1}-v_{i}}\,.$ (4.1)
Here we define $v_{n+1}=v_{1}$. $L$ and $\widehat{L}$ differ in the last interaction term $\lambda\mathrm{e}^{v_{n}-v_{1}}$ corresponding to the extra root in the affine Lie algebra. The constraint $\sum_{i}v_{i}=0$ is consistent because the interactions depend only on the differences. The equations of motion are
$\Delta v_{i}=\lambda\big(e^{v_{i+1}-v_{i}}-e^{v_{i}-v_{i-1}}\big)\,$, (4.2)
taking $v_{i}=v_{i+n}$ in the extended case, but $\exp v_{-1}=\exp v_{n+1}=0$ in the non-extended case. [^16] The theories are often written in terms of the alternative variables $u_{i}=v_{i}-v_{i+1}$, which satisfy $\Delta u_{i}+\lambda\big(2e^{u_{i}}-e^{u_{i-1}}-e^{u_{i+1}}\big)=0$ in the extended case. The case $n=2$ contains the sinh/sine Gordon equations.
These systems are shown to be a reduction of Hitchin’s equations in [MasonWoodhouse199605, §6.2]. This is done by imposing equivariance under a continuous symmetry in the non-extended case and a discrete symmetry in the extended case. In that work they are understood as additional rotational symmetries in the reduction of the 4d self-dual Yang–Mills equations that act on the 2-plane of translational symmetry. It naturally acts also on the spectral parameter, $\zeta\rightarrow\mathrm{e}^{\mathrm{i}\theta}\zeta$ and hence in our context, this symmetry naturally acts on our 4d-Chern–Simons theory too. The symmetries are lifted to the bundle so as to act by the matrices
@@ 4.2 Reduction to Two Dimensions
We now show how these additional symmetries act on our reduction of 4d Chern–Simons to 2d actions so as to land on actions for Toda field theory. Following through the reduction of 4dCS to the Hitchin system we introduce $f$ satisfying (2.19)
@@
Thus $h=\,\mathrm{diag}\{h_{1},\ldots,h_{r}\}$ is diagonal and since it also has unit determinant, we can write
@@
$\psi=\mathrm{i}\begin{pmatrix}0&\psi_{1}&0&\dots&0\\ 0&0&\psi_{2}&\dots&0\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ 0&0&0&\dots&\psi_{r-1}\\ \psi_{r}&0&0&\dots&0\\ \end{pmatrix}\,,\hskip 18.49988pt\widetilde{\psi}=\mathrm{i}\begin{pmatrix}0&0&\dots&0&\widetilde{\psi}_{r}\\ \widetilde{\psi}_{1}&0&\dots&0&0\\ 0&\widetilde{\psi}_{2}&\dots&0&0\\ \vdots&\vdots&\ddots&\vdots&\vdots\\ 0&0&\dots&\widetilde{\psi}_{r-1}&0\\ \end{pmatrix}\,$. (4.9)
@@
$\mathrm{i}\mathop{\text{\Large$\int_{\text{\normalsize$\scriptstyle\Sigma$}}$}}\nolimits\sum_{i=1}^{n}\Big(\partial v_{i}\wedge\bar{\partial}v_{i}+2e^{v_{i}-v_{i+1}}\partial\psi_{i}\wedge\bar{\partial}\widetilde{\psi}_{i}\Big)\,.$ (4.10)
the residual gauge freedoms (2.36) and (2.34), with a slight abuse of notation, replacing $g_{0},g_{\infty}$ by $\mathrm{diag}\{\mathrm{e}^{g_{0i}}\}$ and $\mathrm{diag}\{\mathrm{e}^{g_{\infty i}}\}$ respectively, become
$\delta v_{i}=g_{0i}(w)+g_{\infty i}(\widetilde{w})\,,\hskip 18.49988pt\delta(\psi_{i},\widetilde{\psi}_{i})=(\widetilde{\chi}_{i}(\bar{w}),\chi_{i}(w))\,,\hskip 18.49988pt\sum_{i}g_{0i}=0=\sum_{i}g_{\infty i}\,.$ (4.11)
@@
$\bar{\partial}(e^{v_{i}-v_{i+1}}\partial\psi_{i})=\partial(e^{v_{i}-v_{i+1}}\bar{\partial}\widetilde{\psi}_{i})=0\,.$ (4.12)
@@
$\partial_{w}\psi_{i}=2F_{i}(w)e^{v_{i+1}-v_{i}}\,,\hskip 18.49988pt\partial_{\bar{w}}\widetilde{\psi}_{i}=2\widetilde{F}_{i}(\bar{w})e^{v_{i+1}-v_{i}}\,.$ (4.13)
The $(g_{i},\widetilde{g}_{i})$ can be chosen to make all the $F_{i}=F$ and $\widetilde{F}_{i}=\widetilde{F}$ for some $F(w),\widetilde{F}(\bar{w})$ in the extended case, and in the non-extended case we have enough freedom to reduce to $F=\mathrm{i}=\widetilde{F}$. We can now plug these back into the $v_{i}$ equation of motion to get
$\Delta v_{i}=F(w)\widetilde{F}(\bar{w})(e^{v_{i+1}-v_{i}}-e^{v_{i}-v_{i-1}})$ (4.14)
Thus, we arrive at a slight generalization of the Toda field equation, which specialises to standard form upon fixing $F=1=\widetilde{F}$ [Mikhailov197910, LeznovSaveliev197909, MikhailovOlshanetskyPerelomov198007, OliveTurok198208]. This can always be achieved in the non-extended case. In the extended case we obtain the standard field equation with the coordinate choice
$\mathrm{d}w^{\prime}=F(w)\mathrm{d}w\,.$ (4.15)
This dichotomy between affine and non-affine Toda reflects the conformal invariance of the non-affine case in that the conformal transformation $w\rightarrow w^{\prime}$ is naturally absorbed into the gauge freedom (4.11) whereas in the affine case a coordinate is naturally singled out by (4.15).

### v2

@@ 3.1 $\mathbb{CP}^{1}$-Family of 4d CS Theories and the Corresponding 2d Actions
Since there is a whole $\mathbb{CP}^{1}$ family of hyperkähler structures on $\mathscr{M}_{\text{H}}(\Sigma,G)$, it is expected from our computations in §2.4 that there exists a family of 4d Chern–Simons theories that produces this hyperkähler family of symplectic structures. For this matter, we propose the following $\mathbb{CP}^{1}$-family of meromorphic $(1,0)$-form on $C$
@@
giving the meromorphic $(1,0)$-form used in §2.3 that led to the 4d Chern–Simons theory formulation of Hitchin’s equations with symplectic structure $\Omega_{\mathcal{I}}$. We will see that the $\zeta$ above will be the twistor parameter $\zeta$ of the Hitchin hyperkähler structure, parametrizing the symplectic structures in (1.4). Here we will work with complexified fields; Euclidean reality conditions can be imposed in direct analogy with section 2.3.
@@
These boundary conditions coalesce to those in (2.16) as $\zeta\to 0$. We will find it useful to write $\widehat{\zeta}=-1/\bar{\zeta}$ for the stereographic coordinate of the point antipodal to $\zeta$.
@@
To determine the resulting 2d action, we use the same procedure as explained in the proof of Lemma 2.2, which yields $\frac{\mathrm{i}}{4\pi}S_{\mathrm{H},\zeta}$ where
@@
$-2\mathrm{i}\zeta\mathop{\text{\Large$\int_{\text{\normalsize$\scriptstyle\Sigma$}}$}}\nolimits\mathrm{tr}(k^{-1}\partial k\wedge\,\bar{\partial}\widetilde{\eta})-2\mathrm{i}\bar{\zeta}\int_{\Sigma}\mathrm{tr}(\partial\eta\wedge\bar{\partial}kk^{-1})\,$,
where $S_{\text{WZW}}[k]$ is the WZW action given in equation (2.8). It is straightforward to construct a version of this action depending on gauge covariant data by substituting
$(k,\eta,\widetilde{\eta})=(f_{\widehat{\zeta}}^{-1}f_{\zeta},f_{\hat{\zeta}}^{-1}\boldsymbol{\eta}f_{\hat{\zeta}},f_{\zeta}^{-1}\widetilde{\boldsymbol{\eta}}f_{\zeta})\,.$ (3.11)
The result is a $\zeta$-dependent counterpart of (LABEL:eq:unitary-action).
@@
By taking linear combinations of the above, it’s straightforward to see that
@@
The 4d Chern–Simons reduction to Harmonic maps or principal chiral models are treated in detail in [CostelloYamazaki201908]. We first remark that this also uses an $\omega$ of the form (3.1), but the boundary conditions used are to instead impose $\mathsf{A}=0$ at $\zeta$ and $1/\bar{\zeta}$ instead of (3.4). The Wilson line from $\zeta$ to $1/\bar{\zeta}$ then gives the (twisted) harmonic maps of [Donaldson198707, Corlette198801]. The Lax pair thereby obtained are gauge equivalent to ours, but the twisted harmonic map is realized in a different gauge from our $k$. ∎
@@ 3.2 Hyperkähler Family of Symplectic Structures
In §3.1, we derive a family of 2d actions (3.10) parametrized by $\zeta\in\mathbb{CP}^{1}$ whose equations of motion, for each value of $\zeta$, are Hitchin’s equations (1.3). However, our aim was to find the relationship between $\zeta$ and the twistor parameter of the Hitchin moduli space $\zeta$, which determines the family of symplectic structures (1.4). To this aim, we proceed as we did in §2.4 and compute the symplectic structure that 4d Chern–Simons theory with the choice (3.1) associates to $\Sigma\times S^{1}$, as discussed in §2.4.
@@
By varying $\zeta$ over $\mathbb{CP}^{1}$, (3.19) generates the whole $\mathbb{CP}^{1}$-family of symplectic structures associated to the hyperkähler structure on $\mathscr{M}_{\text{H}}(\Sigma,G)$. Our $\mathbb{CP}^{1}$-family of 4d Chern–Simons theory using $\omega(\zeta,\bar{\zeta})$ of (3.1) generates the $\mathbb{CP}^{1}$-family of 2d actions (3.10) giving the symplectic structures $\Omega_{\mathcal{J}_{\zeta}}$ on $\mathscr{M}_{\text{H}}(\Sigma,G)$, these being the Kähler 2-forms for the complex structures $\mathcal{J}_{\zeta}$ on the space of solutions.
@@ 4 Affine Toda Theory
The complexified Hitchin equations are, up to gauge, the generic reduction of the self-dual Yang–Mills from 4d to 2d along a non-degenerate pair of translations. Such reductions are described in [MasonWoodhouse199605, §6.2] and include a number of familiar 2d integrable systems, including harmonic maps, Toda field theory, and the sine-Gordon equation. These are distinguished as reductions by gauge choices and choices of ‘constants of integration’ that naturally arise in the reductions. When these models are reformulated as 4d Chern–Simons theories, these choices are encoded in the choice of $\omega$ and the boundary conditions on $\mathsf{A}$. It turns out that the harmonic map reductions require different such choices – these are already well-understood in [CostelloYamazaki201908] and require $\omega$ to have two double poles and two single zeroes with $\mathsf{A}$ vanishing at the double poles and allowed to have poles at the zeroes.
However, here we see that Toda field theories fit into the choices made for the Hitchin equations in §2 and can be understood as arising from the imposition of further symmetries. This is a continuous symmetry for Toda field theory (here referring to Toda field theory defined using a Cartan matrix of finite type), but for affine Toda field theory (defined using a Cartan matrix of affine type), we instead impose a discrete symmetry. Toda field theory is a CFT [Braaten:1983pz, Gervais:1983ry, Mansfield:1982sq], whereas affine Toda field theory is integrable but not conformal [Mikhailov197910, LeznovSaveliev197909, MikhailovOlshanetskyPerelomov198007, OliveTurok198208].
Remark 4.1.
Let us briefly survey a number of connections between Chern–Simons theory and Toda field theories appearing in the literature. Toda field theory can be realised as constrained, or equivalently as a gauged, WZW model [Forgacs:1989ac, Balog:1990mu]. This presentation has been exploited to engineer complex Toda theory in 3d Chern–Simons theory with Nahm pole boundary conditions [Cordova:2016cmu], and a lift to a ‘doubled’ 4d Chern–Simons theory was identified in [Stedman202009]. A 4d Chern–Simons realisation of affine Toda theory appears in [FukushimaSakamotoYoshida202112] using the same meromorphic $1$-form $\omega$, but with order defects. We anticipate that our treatment can be connected to theirs by integrating out local degrees of freedom living on the defects. Homogeneous Sine-Gordon models, including in particular complex sine-Gordon, were engineered in [ColeCullinanHoareLiniadoThompson202407] by reducing a gauged holomorphic Chern–Simons theory on twistor space to 4d Chern–Simons. Finally, we note that 4d Chern–Simons with Nahm pole boundary conditions in the topological direction was related to a 3d variant of Toda theory in [AshwinkumarPngTan202008]. ∎
@@ 4.1 Toda Field Theories
Toda theories can be defined for any Kac–Moody algebra, or equivalently for any generalised Cartan matrix, but for simplicity we restrict to the case of $\mathfrak{su}(n)$ for the $A_{n-1}$ series and its untwisted affine counterpart $\widehat{\mathfrak{su}}(n)$. Both cases can be treated together if we introduce variables $v_{1},\ldots,v_{n}$ with $\sum_{i}v_{i}=0$. We then have Lagrangians
$L=\sum_{i=1}^{n}2\partial_{w}v_{i}\partial_{\bar{w}}v_{i}+\lambda\sum_{i=1}^{n-1}\mathrm{e}^{v_{i}-v_{i+1}}\,,\hskip 18.49988pt\widehat{L}=\sum_{i=1}^{n}\big(2\partial_{w}v_{i}\partial_{\bar{w}}v_{i}+\lambda\mathrm{e}^{v_{i}-v_{i+1}}\big)\,.$ (4.1)
In the affine case we take the indices to be cyclic $v_{i}=v_{n+i}$. $L$ and $\widehat{L}$ differ in the last interaction term $\lambda\mathrm{e}^{v_{n}-v_{1}}$ corresponding to the extra root in the affine Lie algebra. The constraint $\sum_{i}v_{i}=0$ is consistent because the interactions depend only on the differences. The equations of motion are
$\Delta v_{i}=\lambda\big(e^{v_{i}-v_{i+1}}-e^{v_{i-1}-v_{i}}\big)\,$, (4.2)
where we understand $\exp v_{0}=\exp v_{n+1}=0$ in the finite case. [^16] Both finite and affine Toda theories are often expressed in terms of alternative variables $t_{i}$ obeying $v_{i}=t_{i}-t_{i-1}$. In the finite case $t_{0}=t_{n}=0$ whereas in the affine case $t_{0}=t_{n}$ and $\sum_{i}t_{i}=0$. They satisfy $\Delta t_{i}=\lambda\exp\sum_{j}K_{ij}t_{j}$ for $K$ the Cartan matrix of $\mathfrak{su}(n)$ or $\widehat{\mathfrak{su}}(n)$ as appropriate. The case $n=2$ contains the sinh/sine-Gordon equations.
These systems are shown to be a reduction of Hitchin’s equations in [MasonWoodhouse199605, §6.2]. This is done by imposing equivariance under a continuous symmetry in the finite case and a discrete symmetry in the affine case. In that work, they are understood as additional rotational symmetries in the reduction of the 4d self-dual Yang–Mills equations that act on the 2-plane of translational symmetry. It naturally acts also on the spectral parameter, $\zeta\rightarrow\mathrm{e}^{\mathrm{i}\theta}\zeta$, and hence in our context, this symmetry naturally acts on our 4d-Chern–Simons theory too. The symmetries are lifted to the bundle so as to act by the matrices
@@ 4.2 Reduction to Two Dimensions
We now show how these additional symmetries act on our reduction of 4d Chern–Simons to 2d actions so as to land on actions for affine Toda field theory. Following through the reduction of 4dCS to the Hitchin system we introduce $f$ satisfying (2.19)
@@
Thus $h=\,\mathrm{diag}\{h_{1},\ldots,h_{n}\}$ is diagonal, and since it also has unit determinant, we can write
@@
$\psi=\begin{pmatrix}0&\psi_{1}&0&\dots&0\\ 0&0&\psi_{2}&\dots&0\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ 0&0&0&\dots&\psi_{n-1}\\ \psi_{n}&0&0&\dots&0\\ \end{pmatrix}\,,\hskip 18.49988pt\widetilde{\psi}=\begin{pmatrix}0&0&\dots&0&\widetilde{\psi}_{n}\\ \widetilde{\psi}_{1}&0&\dots&0&0\\ 0&\widetilde{\psi}_{2}&\dots&0&0\\ \vdots&\vdots&\ddots&\vdots&\vdots\\ 0&0&\dots&\widetilde{\psi}_{n-1}&0\\ \end{pmatrix}\,$. (4.9)
@@
$\mathrm{i}\mathop{\text{\Large$\int_{\text{\normalsize$\scriptstyle\Sigma$}}$}}\nolimits\sum_{i=1}^{n}\Big(\partial v_{i}\wedge\bar{\partial}v_{i}-2e^{v_{i+1}-v_{i}}\partial\psi_{i}\wedge\bar{\partial}\widetilde{\psi}_{i}\Big)\,.$ (4.10)
Slightly abusing notation by writing $\mathrm{diag}\{\mathrm{e}^{g_{0i}}\}$ and $\mathrm{diag}\{\mathrm{e}^{g_{\infty i}}\}$ for $g_{0}$ and $g_{\infty}$ respectively, the residual gauge freedoms (LABEL:psi-gauge) and (2.36) become
$v_{i}\mapsto v_{i}+g_{0i}(w)-g_{\infty i}(\bar{w})\,,\hskip 18.49988pt\sum_{i}g_{0i}=\sum_{i}g_{\infty i}=0\,$, (4.11)
$(\widetilde{\psi}_{i},\psi_{i})\mapsto\big(e^{g_{0i}(w)-g_{0(i+1)}(w)}\widetilde{\psi}_{i},e^{g_{\infty(i+1)}(\bar{w})-g_{\infty i}(\bar{w})}\psi_{i}\big)+(\chi_{i}(w),\widetilde{\chi}_{i}(\bar{w}))\,.$
@@
$\bar{\partial}(e^{v_{i+1}-v_{i}}\partial\psi_{i})=\partial(e^{v_{i+1}-v_{i}}\bar{\partial}\widetilde{\psi}_{i})=0\,.$ (4.12)
@@
$\partial_{w}\psi_{i}=\frac{1}{2}F_{i}(w)e^{v_{i}-v_{i+1}}\,,\hskip 18.49988pt\partial_{\bar{w}}\widetilde{\psi}_{i}=\frac{1}{2}\widetilde{F}_{i}(\bar{w})e^{v_{i}-v_{i+1}}\,.$ (4.13)
The $(g_{i},\widetilde{g}_{i})$ can be chosen to make all the $F_{i}=F$ and $\widetilde{F}_{i}=\widetilde{F}$ for some $F(w),\widetilde{F}(\bar{w})$ in the affine case, and in the finite case we have enough freedom to reduce to $F=\widetilde{F}=1$. We can now plug these back into the $v_{i}$ equation of motion to get
$\Delta v_{i}=F(w)\widetilde{F}(\bar{w})(e^{v_{i}-v_{i+1}}-e^{v_{i-1}-v_{i}})\,.$ (4.14)
In the finite case with $F=\widetilde{F}=1$ these are the Toda field equations. For the affine Kac-Moody algebra $\widehat{\mathfrak{su}}(n)$ they are a (subset of) the equations of motion of conformal affine Toda theory (CAT) [BabelonBonora199007, AratynFerreiraGomesZimerman199107, Constantinidis:1992hs]. [^17] The equations of motion of CAT take the form $\Delta t_{i}=e^{\eta}e^{\sum_{j}K_{ij}t_{j}}$ for $K$ the Cartan matrix of $\widehat{\mathfrak{su}}(n)$ and $\eta$ an additional field obeying $\Delta\eta=0$. We can identify $F(w)\widetilde{F}(\bar{w})=e^{\eta}$. The constraint $\sum_{i}t_{i}=0$ is also not imposed, but since the interactions depend only on differences, this extra centre of mass variable does not enter the $v_{i}=t_{i}-t_{i-1}$ equations of motion. We are grateful to Luiz Agostinho Ferreira for alerting us to the connection to CAT. Making the coordinate choice
$\mathrm{d}w^{\prime}=F(w)\mathrm{d}w$ (4.15)
recovers the standard field equations of affine Toda theory. This dichotomy between finite and affine Toda reflects the conformal invariance of the former in that the conformal transformation $w\rightarrow w^{\prime}$ is naturally absorbed into the gauge freedom (4.11) whereas in the latter a coordinate is naturally singled out by (4.15).
Remark 4.2.
The action (4.10) does not match the familiar expressions given in equation (4.1). Indeed, it depends on an extra field $\psi,\widetilde{\psi}$ which can be eliminated on-shell. These extra degrees of freedom play a crucial role in the Alday-Maldacena formulation of minimal surfaces, but it would nevertheless be desirable to recover the Lagrangians (4.1) directly. We expect that in order to do so, conformal symmetry must be explicitly broken through the boundary conditions. ∎